OpenStudy (tbnrfrags):

What is the remainder when 4^31 is divided by 10? A:2 B:4 C:8 D:6 E:0

OpenStudy (carolina4567):

B 4

OpenStudy (tmak11):

well 4^31 is 4611686018427387904 and dividing by....hes correct^^^^^

OpenStudy (carolina4567):

Im a SHE :)

OpenStudy (tmak11):

oops im sorry o have a friend named caroline so im so used to saying he

OpenStudy (tbnrfrags):

ty<3

OpenStudy (anonymous):

he answer is b

OpenStudy (tbnrfrags):

<3

OpenStudy (carolina4567):

4^31 = 4*16^15 = 4*6 = 4

OpenStudy (mathmath333):

sry \(\large\tt \color{black}{\dfrac{4^{31}}{10}}\) \(\large\tt \color{black}{\dfrac{=4\times6^{15}}{10}}\) \(\large\tt \color{black}{=\dfrac{4\times6}{10}}\)

ganeshie8 (ganeshie8):

nice :)

OpenStudy (mathmath333):

thks

OpenStudy (anonymous):

b:4